OpenStudy (kainui):

How to multiply surfaces as matrices intro:

OpenStudy (dan815):

btw, can this be thought of as a version of convolutions then

OpenStudy (kainui):

We can choose a matrix to be any place on a surface, it doesn't matter where. Just as long as they're evenly placed it will work out. So we are going to define our matrix \[\LARGE A=\left[\begin{matrix}-1 & -2 \\ 2& 8\end{matrix}\right]\]\[\LARGE A(x,y)=(-y)^xx\] So here I'm saying we're taking samples at steps of 1 and going from 1 to 2 on the x and y axes. We can sample the same function in different areas different ways and get rectangular matrices or whatever, and it's quite easy to show that \[\LARGE A^T(x,y)=A(y,x)\] Now for multiplication: \[\LARGE \sum_{n=1}^2A(x,n)B(n,y)\] We are simply going over the 2x2 matrix over the bounds. If B is really A on the same surface, we should scale it so that at n=1 and n=2 we hit the right spots. This is just like how with matrices we have: \[\LARGE A_{ij}B_{jk}=C_{ik}\] So let's let B be the identity function, \[\LARGE \delta (x,y)\] when x=y it is 1, otherwise it is zero. \[\LARGE A=AI=\sum_{n=1}^2A(x,n) \delta(n,y)\]\[\LARGE A(x,1) \delta(1,y)+A(x,2) \delta(2,y)=A(x,y)\] How do I know this statement is true? Just evaluate it at y=1 or y=2 and see that you will drop off all the other terms. Whether you left or right multiply simply sorts it out into its rows or columns.

OpenStudy (kainui):

Notice though, all of these functions are secretly not discrete though. So the product AB is generally another surface that we can sample from. And we also have the ability to sorta just sample infinitely between points and turn this into an integral. Or collapse our matrices into single vectors and it's just the regular old dot product we had with integrals in the past.

OpenStudy (dan815):

does it even matter how many points u can select for A, or if it has to be symmetric or anything, is it fully dependant on just the functions and knroicker

OpenStudy (kainui):

You can select as many points as you like. It just has to obey the regular matrix multiplication rules that the inside parts of their indices line up, you can multiply a 3x7 times a 7x5 matrix just as easily and your summation will run from 1 to 7.

OpenStudy (dan815):

oh okay so like u did x=1,y=1 then x=1,y=2 2 1 2 2

OpenStudy (dan815):

ok i can see why the transform is true now

OpenStudy (kainui):

If you recall watching this video, these Fourier Matrices are quite nicely expressed as functions. https://www.youtube.com/watch?v=umt6BB1nJ4w&list=PL49CF3715CB9EF31D&index=26 and their inverse is quite easy to compute and see that by following the rules I wrote above will give us the identity surface as well.

OpenStudy (kainui):

Every surface is the sum of a symmetric and skew symmetric matrix \[\LARGE A(x,y)=B(x,y)+C(x,y)\] This isn't very surprising but hey it's useful since it's not very clear how to actually represent an arbitrary matrix with functions.\[\LARGE B(x,y)=\frac{A(x,y)+A(y,x)}{2}\] \[\LARGE C(x,y)=\frac{A(x,y)-A(y,x)}{2}\]

OpenStudy (kainui):

But really just like Fourier series it seems like it shouldn't be too big of a deal.

OpenStudy (dan815):

ok i got what u said so far

OpenStudy (dan815):

what do u want me to watch in this lecture there no fseries

OpenStudy (kainui):

https://www.youtube.com/watch?v=M0Sa8fLOajA&list=PL49CF3715CB9EF31D#t=1148 There, sorry I think I sent you the wrong thing.

OpenStudy (kainui):

Essentially here is the formula in my scheme\[\LARGE (F_n)_{ab}=e^{i \frac{2\pi}{n}ab}\] and like he says earlier, we're starting this at k=0 and go up to k=n-1 as the dimensions of the matrix, so we're creating an nxn matrix this way.

OpenStudy (dan815):

wait im rewatching from hemetian matrices part in this lecture

OpenStudy (dan815):

he just started F_n

OpenStudy (kainui):

The inverse is just the hermitian of this matrix, which is quite obvious to see the conjugate transpose is really just 1/Fn really convenient haha.

OpenStudy (dan815):

thats cool u can also tell those columns had to be orthogonal because of that inner product of and sin and cos functions showing up in the inner products

OpenStudy (dan815):

and u know that since its different roots m=/=n in the sin(mx)*cos(nx) or cos(mx)*cos(nx) represenation of their innerproducts

OpenStudy (dan815):

u know m =/=n because theyre all for a different set of roots along each column or row

OpenStudy (kainui):

So like multiplication wise we have: \[\LARGE \sum_{k=0}^{n-1}e^{i \frac{2\pi}{n}xk}e^{-i \frac{2\pi}{n}ky}=\delta(x,y)\]

OpenStudy (dan815):

right okay thats what u mean u had complex representaion of this identity

OpenStudy (dan815):

hmm id have to see this thru tho

OpenStudy (dan815):

ok wait i see it

OpenStudy (dan815):

hmmm wait

OpenStudy (dan815):

how is it zero for x=/= y

OpenStudy (dan815):

F_n^H*F_n=I

OpenStudy (kainui):

Try simple cases like n=2, n=3, n=4.

OpenStudy (kainui):

actually there's a 1/sqrt(n) factor out front of each of these matrices I forgot

OpenStudy (dan815):

ya ya np

OpenStudy (kainui):

The only thing is it will give you instead of the identity will be a scalar matrix with that n value.

OpenStudy (dan815):

ur function is the identiy if x ,y we chose are integers

OpenStudy (dan815):

are u tryna now say we can continue this to non discrete

OpenStudy (kainui):

sure thing d000d

OpenStudy (dan815):

lol

OpenStudy (dan815):

tell me

OpenStudy (dan815):

i think we are starting fourier transforms again soon, gotta revisit this then

OpenStudy (dan815):

now imma start my exam-.-

OpenStudy (kainui):

ok have fun I'm playing around right now trying to figure stuff out. Keep in mind this means we can also do stuff like tensors and have A(x,y,z) and then by permuting the variables we have higher order transposes.

OpenStudy (kainui):

Here's a fun question, what's the transpose of a matrix in polar form?

OpenStudy (dan815):

the identiy wud be equivalent to e^itehta

OpenStudy (dan815):

well e^i*2pi

OpenStudy (dan815):

ud have to get a matrix in its polar orthonormal form and maybe the transfmore of that wud be its inverse in polar too

OpenStudy (dan815):

transpose*

OpenStudy (dan815):

hmm theres gotta be some way to extract a general transpose in polar form from these properties

OpenStudy (kainui):

\[\LARGE A^T(r(x,y),\theta(x,y))=A(r(y,x),\theta(y,x))\] \[\LARGE r=\sqrt{x^2+y^2}=r^T\]\\[\LARGE \theta^T=\tan^{-1} \frac{x}{y}=\tan^{-1}\cot \theta\]

OpenStudy (dan815):

oh cool!

OpenStudy (dan815):

A^T=A(r^T,theta^T) O_O

OpenStudy (dan815):

what can we do with this

OpenStudy (dan815):

feel like we are just saying something with matrices what we can say anyway by switching axis

OpenStudy (kainui):

Yeah that's all we're doing really. It's just sorta a lame playing around of stuff to see what it looks like for fun.

OpenStudy (dan815):

ahh ok vcarry on http://dmpeli.math.mcmaster.ca/TeachProjects/Math3C03_14/home4.pdf

OpenStudy (dan815):

see if anything peaks ur interest like the last question perhaps -.-

OpenStudy (dan815):

lmao the most random first comment here